# Continual Release Moment Estimation with Differential Privacy

### v1

@@ C.1 Bounds on the expected approximation error for PP with non-trivial factorization
$\mathbb{E}_{Z_{1},Z_{2}}\|V(X\operatorname{\!\bullet\!}X+\lambda^{-1/2}Z_{2})-(V(X+Z_{1})\operatorname{\!\bullet\!}V(X+Z_{1}))-V(X\operatorname{\!\bullet\!}X)+(VX)\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}$
@@
$\quad=\mathbb{E}_{Z_{2}}\|\lambda^{-1/2}VZ_{2}\|_{\text{F}}^{2}+2\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}+\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}+2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\!\bullet\!}(VX),(VX)\operatorname{\!\bullet\!}(VZ_{1})\rangle$
$\quad=c_{d}\sigma^{2}d^{2}\|V\|_{\text{F}}^{2}+2{\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}}+{\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}}+2{\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\!\bullet\!}(VX),(VX)\operatorname{\!\bullet\!}(VZ_{1})\rangle}$
Using Proposition C.2 and Proposition C.3, we therefore have
@@
$\mathbb{E}_{Z_{1},Z_{2}}\|V(X\operatorname{\!\bullet\!}X+Z_{2})-(V(X+Z_{1})\operatorname{\!\bullet\!}V(X+Z_{1}))-V(X\operatorname{\!\bullet\!}X)+(VX)\operatorname{\!\bullet\!}(VX)+\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}$
@@
$\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}=\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}=d\sigma^{4}H_{n,2}.$
@@
$-2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}),\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\rangle+\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}$ $=d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}=d\sigma^{4}H_{n,2}.$
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}\|_{\text{F}}^{2}$ $=\sigma^{2}(c_{d}d^{2}+2d+2)\sum\limits_{k=1}^{n}\frac{1}{k}+d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}$
@@
$\begin{split}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}=&\mathbb{E}_{Z_{1}}\|V((X+Z_{1})\operatorname{\!\bullet\!}(X+Z_{1}))-(V(X+Z_{1})\operatorname{\!\bullet\!}V(X+Z_{1}))-V(X\operatorname{\!\bullet\!}X)+(VX)\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}\\ &\quad=\mathbb{E}_{Z_{1}}\|V(Z_{1}\operatorname{\!\bullet\!}Z_{1})+V(Z_{1}\operatorname{\!\bullet\!}X)+V(X\operatorname{\!\bullet\!}Z_{1})-(VZ_{1})\operatorname{\!\bullet\!}(VX)-(VX)\operatorname{\!\bullet\!}(VZ_{1})-(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}))\|_{\text{F}}^{2}\\ &\quad=\underbrace{\mathbb{E}_{Z_{1}}\|V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\|_{\text{F}}^{2}}_{A_{1}}-2\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}Z_{1}),(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\rangle}_{A_{2}}+2\underbrace{\mathbb{E}_{Z_{1}}\|V(Z_{1}\operatorname{\!\bullet\!}X)\|_{\text{F}}^{2}}_{A_{3}}\\ &\hskip 56.9055pt+2\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}X),V(X\operatorname{\!\bullet\!}Z_{1})\rangle}_{A_{4}}-4\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}X),(VX)\operatorname{\!\bullet\!}(VZ_{1})\rangle}_{A_{5}}\\ &\hskip 56.9055pt-4\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}X),(VZ_{1})\operatorname{\!\bullet\!}(VX)\rangle}_{A_{6}}+2\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}+\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}\\ &\hskip 56.9055pt+2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\!\bullet\!}(VX),(VX)\operatorname{\!\bullet\!}(VZ_{1})\rangle\end{split}$ (63)
@@
$\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}+d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}\right)^{2}-2d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{3}$
@@
$\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}X_{t,j}^{2}-4(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}X_{t,j}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}X_{t,j}\right)$
@@
$\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-2d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}-2d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}$
@@
$=d(d-1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-d^{2}\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2(d+1)\sigma^{2}\underbrace{\sum\limits_{k=1}^{n}\frac{1}{k^{2}}\left[\sum\limits_{t=1}^{k}\langle X_{t,:},X_{t,:}\rangle-\frac{1}{k}\sum\limits_{t_{1},t_{2}}^{k}\langle X_{t_{1},:},X_{t_{1},:}\rangle\right]}_{T_{n}(X)}$
@@
$T_{n}:=\sup_{X\in\mathcal{X}}T_{n}(X)\sup_{X\in\mathcal{X}}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}\left[\sum\limits_{t=1}^{k}\langle X_{t,:},X_{t,:}\rangle-\frac{1}{k}\underbrace{\sum\limits_{t_{1},t_{2}}^{k}\langle X_{t_{1},:},X_{t_{2},:}}_{=\left\|\sum_{t}X_{t}\right\|^{2}_{2}\geq 0}\rangle\right].$
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $\leq d(d-1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-d^{2}\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k}$
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $\geq d(d-1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-d^{2}\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k}-2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k^{3}},$
@@
$\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}$ $=\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}$
$\|\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\|_{\text{F}}^{2}$ $=\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}\right)^{2}=dn\sigma^{4}$
$\langle\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}),\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\rangle_{\text{F}}$ $=d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}\right)=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}$
@@
$\|\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\|_{\text{F}}^{2}+\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}-2\langle\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}),\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\rangle$
@@
Combining everything together we obtain:
$\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ $=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}-2d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{3}$
$\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}X_{t,j}^{2}-4(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}X_{t,j}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}X_{t,j}\right)$
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ $\leq d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}+2(d+1)\sigma^{2}H_{n,1}\quad\text{and}$
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ $\geq d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}+2(d+1)\sigma^{2}H_{n,1}-2(d+1)\sigma^{2}H_{n,3},$
@@
$f(x_{i},y_{i})$ $:=\left(\sqrt{x_{i}^{2}+x_{j}^{2}}+\sqrt{y_{i}^{2}+y_{j}^{2}}\right)^{2}+\lambda\left(x_{i}^{2}+x_{j}^{2}-y_{i}^{2}-y_{j}^{2}\right)^{2}$
$g(x_{i},y_{i})$ $:=(x_{i}-y_{i})^{2}+(x_{j}-y_{j})^{2}+\lambda(x_{i}^{2}-y_{i}^{2})^{2}+\lambda(x_{j}^{2}-y_{j}^{2})^{2}.$
@@
Lemma C.4 (Expected Second Moment Error with PP).
@@
$=2d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))+4\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\cdot\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{T})$
$\quad+\underbrace{d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})}_{\text{bias}}$
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}((X+C_{1}^{-1}Z_{1})\circ(X+C_{1}^{-1}Z_{1}))-A_{2}(X\circ X)\|_{\text{F}}^{2}$
@@
$S_{1}$ $=\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}(X\circ C_{1}^{-1}Z_{1})\|_{\text{F}}^{2}=\sup_{X\in\mathcal{X}}\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}(A_{2})_{k,t}X_{t,j}\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t,r}(Z_{1})_{r,j}\right)^{2}$

### v2

@@ D.1 Bounds on the expected approximation error for PP with non-trivial factorization
$\mathbb{E}_{Z_{1},Z_{2}}\|V(X\operatorname{\bullet}X+\lambda^{-1/2}Z_{2})-(V(X+Z_{1})\operatorname{\bullet}V(X+Z_{1}))-V(X\operatorname{\bullet}X)+(VX)\operatorname{\bullet}(VX)\|_{\text{F}}^{2}$ (68)
@@
$\quad=\mathbb{E}_{Z_{2}}\|\lambda^{-1/2}VZ_{2}\|_{\text{F}}^{2}+2\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\bullet}(VX)\|_{\text{F}}^{2}+\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}$
$\qquad\qquad+2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\bullet}(VX),(VX)\operatorname{\bullet}(VZ_{1})\rangle$
$\quad=c_{d}\sigma^{2}d^{2}\|V\|_{\text{F}}^{2}+2{\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\bullet}(VX)\|_{\text{F}}^{2}}+{\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}}$
$\qquad\qquad+2{\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\bullet}(VX),(VX)\operatorname{\bullet}(VZ_{1})\rangle}$
Using Proposition D.3 and Proposition D.4, we therefore have
@@
$\mathbb{E}_{Z_{1},Z_{2}}\|V(X\operatorname{\bullet}X+Z_{2})$ $-(V(X+Z_{1})\operatorname{\bullet}V(X+Z_{1}))-V(X\operatorname{\bullet}X)+(VX)\operatorname{\bullet}(VX)$ (76)
$+\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}$
@@
$\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}=\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}=d\sigma^{4}H_{n,2}.$ (77)
@@
$-2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\bullet}(VZ_{1}),\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1})\rangle+\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}$ $=d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}$ (78)
$=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}=d\sigma^{4}H_{n,2}.$
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}\|_{\text{F}}^{2}$ $=\sigma^{2}(c_{d}d^{2}+2d+2)\sum\limits_{k=1}^{n}\frac{1}{k}+d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}$ (79)
@@
$\begin{split}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}=&\mathbb{E}_{Z_{1}}\|V((X+Z_{1})\operatorname{\bullet}(X+Z_{1}))-(V(X+Z_{1})\operatorname{\bullet}V(X+Z_{1}))-V(X\operatorname{\bullet}X)+(VX)\operatorname{\bullet}(VX)\|_{\text{F}}^{2}\\ &\quad=\underbrace{\mathbb{E}_{Z_{1}}\|V(Z_{1}\operatorname{\bullet}Z_{1})\|_{\text{F}}^{2}}_{A_{1}}-2\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\bullet}Z_{1}),(VZ_{1})\operatorname{\bullet}(VZ_{1})\rangle}_{A_{2}}+2\underbrace{\mathbb{E}_{Z_{1}}\|V(Z_{1}\operatorname{\bullet}X)\|_{\text{F}}^{2}}_{A_{3}}\\ &\hskip 56.9055pt+2\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\bullet}X),V(X\operatorname{\bullet}Z_{1})\rangle}_{A_{4}}-4\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\bullet}X),(VX)\operatorname{\bullet}(VZ_{1})\rangle}_{A_{5}}\\ &\hskip 56.9055pt-4\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\bullet}X),(VZ_{1})\operatorname{\bullet}(VX)\rangle}_{A_{6}}+2\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\bullet}(VX)\|_{\text{F}}^{2}\\ &\hskip 56.9055pt+\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}\\ &\hskip 56.9055pt+2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\bullet}(VX),(VX)\operatorname{\bullet}(VZ_{1})\rangle\end{split}$ (80)
@@
$\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}+d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}\right)^{2}-2d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{3}$ (87)
@@
$\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}X_{t,j}^{2}$
$\quad-4(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}X_{t,j}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}X_{t,j}\right)$
@@
$\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-2d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}-2d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}$ (88)
@@
$=d(d-1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-d^{2}\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}$
$\quad+2(d+1)\sigma^{2}\underbrace{\sum\limits_{k=1}^{n}\frac{1}{k^{2}}\left[\sum\limits_{t=1}^{k}\langle X_{t,:},X_{t,:}\rangle-\frac{1}{k}\sum\limits_{t_{1},t_{2}}^{k}\langle X_{t_{1},:},X_{t_{1},:}\rangle\right]}_{T_{n}(X)}$
@@
$T_{n}:=\sup_{X\in\mathcal{X}}T_{n}(X)\sup_{X\in\mathcal{X}}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}\left[\sum\limits_{t=1}^{k}\langle X_{t,:},X_{t,:}\rangle-\frac{1}{k}\underbrace{\sum\limits_{t_{1},t_{2}}^{k}\langle X_{t_{1},:},X_{t_{2},:}}_{=\left\|\sum_{t}X_{t}\right\|^{2}_{2}\geq 0}\rangle\right].$ (89)
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $\leq d(d-1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-d^{2}\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k}$ (91)
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $\geq d(d-1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-d^{2}\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k}-2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k^{3}},$ (92)
@@
$\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}$ $=\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}$ (94)
$\|\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\bullet}Z_{1})\|_{\text{F}}^{2}$ $=\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}\right)^{2}=dn\sigma^{4}$ (95)
$\langle\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1}),\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\bullet}Z_{1})\rangle_{\text{F}}$ $=d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}\right)=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}$ (96)
@@
$\|\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\bullet}Z_{1})\|_{\text{F}}^{2}+\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}-2\langle\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1}),\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\bullet}Z_{1})\rangle$ (97)
@@
Combining everything together, we obtain:
$\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ $=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}-2d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{3}$ (98)
$\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}X_{t,j}^{2}$
$\quad-4(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}X_{t,j}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}X_{t,j}\right)$
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ $\leq d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}+2(d+1)\sigma^{2}H_{n,1}\quad\text{and}$ (99)
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ $\geq d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}+2(d+1)\sigma^{2}H_{n,1}-2(d+1)\sigma^{2}H_{n,3},$ (100)
@@
$f(x_{i},y_{i})$ $:=\left(\sqrt{x_{i}^{2}+x_{j}^{2}}+\sqrt{y_{i}^{2}+y_{j}^{2}}\right)^{2}+\lambda\left(x_{i}^{2}+x_{j}^{2}-y_{i}^{2}-y_{j}^{2}\right)^{2}$ (101)
$g(x_{i},y_{i})$ $:=(x_{i}-y_{i})^{2}+(x_{j}-y_{j})^{2}+\lambda(x_{i}^{2}-y_{i}^{2})^{2}+\lambda(x_{j}^{2}-y_{j}^{2})^{2}.$ (102)
@@
Lemma D.5 (Expected Second Moment Error with PP).
@@
$=2d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))$
$\qquad+4\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\cdot\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{T})$
$\qquad+\underbrace{d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})}_{\text{bias}}$
@@
$\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}((X+C_{1}^{-1}Z_{1})\circ(X+C_{1}^{-1}Z_{1}))-A_{2}(X\circ X)\|_{\text{F}}^{2}$ (129)
@@
$S_{1}$ $=\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}(X\circ C_{1}^{-1}Z_{1})\|_{\text{F}}^{2}=\sup_{X\in\mathcal{X}}\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}(A_{2})_{k,t}X_{t,j}\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t,r}(Z_{1})_{r,j}\right)^{2}$ (130)